# Endomorphism algebras of abelian varieties with large cyclic 2-torsion field over a given field

### v3

@@ 2. Proof of the main theorems
Let $F/K$ be an odd degree abelian extension unramified outside of 2. Then by class field theory, any prime dividing $[F\colon K]$ divides either the class number of $K$, or the multiplicative order of the residue field of a prime above 2 in $K$.
@@
As $E/\mathbb{Q}$ is Galois, the reflex field $E^{*}$ is a subfield of $E$ [Shi98, Prop. 28, pg 62]. Moreover, since $A$ is absolutely simple its CM type is primitive. Proposition 30 on page 65 of [Shi98] applies to tell us the endomorphism field $L$ equals $E^{*}K$. Finally, if $E/\mathbb{Q}$ is an abelian extension, then $E=E^{*}$ by Example (1) on page 63 of [Shi98]. ∎
@@
  • $p\in\{7,11,19,43,67,163\}$ and $A$ is isogenous over $\bar{\mathbb{Q}}$ to the power of an elliptic curve with complex multiplication by $\mathbb{Q}(\sqrt{-p})$.
@@
Clearly, we may apply Theorem 2.4. Suppose first $p$ divides $[L:\mathbb{Q}]$. Then $g\geq 3$ and $A$ is isogenous over $\bar{\mathbb{Q}}$ to the power of an absolutely simple abelian variety with complex multiplication by a proper (non-trivial) subfield $F$ of $\mathbb{Q}(\zeta_{p})$. Thus there is some minimal finite extension $K/\mathbb{Q}$ such that $A_{K}$ contains a simple abelian subvariety $B$ with CM by $F$. The above lemma applies to $B$ and shows $L$, the endomorphism field of $A$, contains $KF$.
Proposition 3.9 of [Goo21] (recalled at the beginning of this section) shows $\operatorname{{\rm Gal}}(L/\mathbb{Q})$ is an extension of $C_{p}$ by $C_{2}^{m}$ for some $m$. In particular, any non-trivial element of even order in $\operatorname{{\rm Gal}}(L/\mathbb{Q})$ has either order 2 or $2p$.
Since $p$ does not divide $[F:\mathbb{Q}]$, the cyclic quotient $\operatorname{{\rm Gal}}(F/\mathbb{Q})$ of $\operatorname{{\rm Gal}}(L/\mathbb{Q})$ has order two. Using $F$ is a CM field only ramified at $p$, we find $F=\mathbb{Q}(\sqrt{-p})$ where $p\equiv 3\mod{4}$. In particular $g$ is odd.
We show $F$ has class number one, which by the Baker–Heegner–Stark Theorem will conclude the proof in this case. As $F/\mathbb{Q}$ is a quadratic extension, $B$ is an elliptic curve. Thus by CM theory [Sil94, II §4 Thm. 4.1, pg 121] $KF$ contains $H$ the Hilbert class field of $F$. In particular, $L$ contains $H$, so $[H:F]$ divides $2^{m}p$ for some $m$. On the other hand, [FG18, Corollary 2.16] tells us every element of $\operatorname{{\rm Gal}}(H/F)$ has order dividing $g$. As $g$ is odd and less than $p$, we find $\operatorname{{\rm Gal}}(H/F)=1$ as claimed.
We now suppose $p$ does not divide $[L:\mathbb{Q}]$ and show $E\neq\mathbb{Q}(\zeta_{p})$. Suppose we had equality, then applying Lemma 2.5 we find $L=\mathbb{Q}(\zeta_{p})$. If $p=3$, then we apply Theorem 2.9 of [Goo21] and Theorem 2.1 to obtain a contradiction. Else, $p\geq 5$ and Proposition 3.9 of [Goo21] provides us with a contradiction, since $\operatorname{{\rm Gal}}(\mathbb{Q}(\zeta_{p})/\mathbb{Q})$ is not an elementary abelian $2$-group. ∎
Let us record some easy corollaries of Theorem 2.9 of [Goo21], Theorems 2.1, 2.6 and Lemma 2.5.
@@
Let $A\colon y^{2}=f(x)$ be an elliptic curve defined over a number field with a real embedding. If $\operatorname{{\rm Gal}}(f)\cong C_{3}$, then $A$ does not have complex multiplication.
@@
Both cases are possible. Indeed, examples of the first case are easily found. For the second case, note the jacobian $J$ of the hyperelliptic curve $y^{2}=x(x^{5}-4x^{4}+2x^{3}+5x^{2}-2x-1)$ has $\operatorname{{\rm End}}_{\mathbb{Q}}(J)=\operatorname{{\rm End}}(J)\cong\mathbb{Z}\left[\frac{1+\sqrt{5}}{2}\right]$ by [Wil00, Proposition 1] and the Galois group of $x^{5}-4x^{4}+2x^{3}+5x^{2}-2x-1$ has order 5.
@@
  • $p\in\{7,11,19,43,67,163\}$ and $A$ is isogenous over $\bar{K}$ to the power of an elliptic curve with complex multiplication by $\mathbb{Q}(\sqrt{-p})$.
@@
Clearly, we may apply Theorem 2.4. Suppose first $p$ divides $[L:K]$. Then $A$ is isogenous over $\bar{K}$ to the power of an absolutely simple abelian variety $B$ with complex multiplication by a proper subfield $F$ of $\mathbb{Q}(\zeta_{p})$. Lemma 2.5 applies to $B$ and shows $L$, the endomorphism field of $A$, contains $F$.
Proposition 3.9 of [Goo21] (recalled at the beginning of this section) shows $\operatorname{{\rm Gal}}(L/K)$ is an extension of $C_{p}$ by $C_{2}^{m}$ for some $m$. In particular, any non-trivial element of even order in $\operatorname{{\rm Gal}}(L/K)$ has either order 2 or $2p$. Since $p$ does not divide $[F:\mathbb{Q}]$, the cyclic quotient $\operatorname{{\rm Gal}}(FK/K)$ of $\operatorname{{\rm Gal}}(L/K)$ has order dividing two.
In turn we deduce $[F:\mathbb{Q}]$ divides $4$ and is equal to 4 only if $F\supsetneq K$. Being a CM field, $[F:\mathbb{Q}]=4$ would imply $F$ contains a real quadratic field. Whence $\operatorname{{\rm Gal}}(F/\mathbb{Q})\cong C_{2}\times C_{2}$ contradicting the fact $F\subseteq\mathbb{Q}(\zeta_{p})$. Thus $[F:\mathbb{Q}]=2$. As $F$ is a CM field only ramified at $p$, we deduce $F=\mathbb{Q}(\sqrt{-p})$ where $p\equiv 3\mod{4}$.
[…]
@@
We now suppose $p$ does not divide $[L:K]$ and show $E\neq\mathbb{Q}(\zeta_{p})$. Suppose we had equality, then by Lemma 2.5, $L=K(\zeta_{p})$. Proposition 3.9 of [Goo21] then implies $\operatorname{{\rm Gal}}(L/K)=\operatorname{{\rm Gal}}(K(\zeta_{p})/K)\cong\operatorname{{\rm Gal}}(\mathbb{Q}(\zeta_{p})/\mathbb{Q}(\zeta_{p})\cap K)$ has order at most $2$. Hence $[\mathbb{Q}(\zeta_{p}):\mathbb{Q}]$ divides $4$, being equal to 4 only if $K$ is contained in $\mathbb{Q}(\zeta_{p})$. But $[\mathbb{Q}(\zeta_{p}):\mathbb{Q}]=4$ implies $p=5$ and $\mathbb{Q}(\zeta_{5})$ does not have an imaginary subfield. Whence $[\mathbb{Q}(\zeta_{p}):\mathbb{Q}]=2$ and $p=3$. This in turn implies $g=1$ which we have ruled out by assumption. ∎
@@
Let $K$ be an imaginary quadratic field of class number divisible by 5. Let $H$ be a degree 5 extension of $K$ contained in the Hilbert class field of $K$. What are the possible endomorphism algebras for abelian surfaces $A/\mathbb{Q}$ with $\mathbb{Q}(A[2])=H$?
@@
By combining Theorem 2.1, Theorem 2.10 of [Goo21] and [Shi98, Prop. 30, pg.65; Example (2), pg.64], we see that either $\operatorname{{\rm End}}^{0}(A)\cong\mathbb{Q}$ or $\mathbb{Q}(\sqrt{d})$ with $d\equiv 5\mod{8}$.
The above question then becomes: given $H$, which $d$ occur?
We finish this section with the following extension of [Goo21, Theorem 3.5].
@@
Let $f\in K[x]$ be a polynomial of odd degree $n$ with Galois group isomorphic to a Frobenius group $G$ of order $n(n-1)$. Let $J_{f}$ be the jacobian associated to the hyperelliptic curve defined by $y^{2}=f(x)$.
Suppose $E={\rm End}^{0}(J_{f})$ is a number field of dimension $s$ over $\mathbb{Q}$. Then $E/\mathbb{Q}$ is Galois with ${\rm Gal}(E/\mathbb{Q})$ isomorphic to a quotient of $H$ the Frobenius complement of $G$.
Furthermore, $L/K$ is an extension of degree $s$ contained in $K(f)$, and as abstract groups, ${\rm Gal}(L/K)\cong{\rm Gal}(E/\mathbb{Q})$. Moreover, if $s=n-1$, then $L=EK$.
@@
This result is [Goo21, Theorem 3.5] with the extra assertion that if $s=n-1$, then $L=EK$. If $s=n-1$, then $J_{f}$ is an absolutely simple abelian variety with CM by $E$. We may therefore apply Lemma 2.5 to find $L=E^{*}K$ and $E^{*}\subseteq E$. Hence $[E^{*}:\mathbb{Q}]\geq[L:K]=s=[E:\mathbb{Q}]$. It follows we have equality $E^{*}=E$. ∎
@@
This field is totally inert at 2 (as predicted by [Goo21, Theorem 2.9]) and unramified outside 13. We note its class number is one, but in line with the theorems presented above, $\mathbb{Q}(f)/K$ is a degree 5 extension ramified only at 2.
@@
Let $B=\mathbb{Q}+i\mathbb{Q}+j\mathbb{Q}+k\mathbb{Q}$ be an indefinite quaternion algebra generated by $i,j$ satisfying $i^{2}=D/m$, $j^{2}=m$, $ij=-ji$ and $k=ij$, where $D$ is a positive squarefree integer and $m|D$. We note that as $B$ is indefinite, $0>-m=\mathrm{n}(j)$, the reduced norm of $j$. As we shall only use the reduced norm, reduced trace and reduced discriminant, we shall refer to these simply as the norm, trace and discriminant respectively.
@@
An order of $B$ is said to be hereditary if it has squarefree (reduced) discriminant (for alternate, equivalent, definitions see [DR04, Voi21]). We shall take $\mathcal{O}$ to be a hereditary order in $B$ of discriminant $D$.
@@
An abelian surface $A/K$ is said to have quaternion multiplication by $\mathcal{O}$ if there is an isomorphism $\iota\colon\mathcal{O}\xrightarrow{\sim}\operatorname{{\rm End}}(A)$. We shall assume there is an element $\mu\in\mathcal{O}$ satisfying $\mu^{2}+D=0$. This is always the case when $\mathcal{O}$ is a maximal order [Voi21, 43.6.6, page 818] and furthermore such an element $\mu$ induces a principal polarisation on $A$ [Voi21, page 821]. Fix such an $A$, $\mu$ and $\iota$.
@@
Following [DR04, Definition 3.3], we say $\chi\in B$ is a twist of $(\mathcal{O},\mu)$ if it lies in both $\mathcal{O}$ and the normaliser $N_{B^{*}}(\mathcal{O})$, has trace zero, $\mathrm{n}(\chi)$ divides $D$, and $\mu\chi=-\chi\mu$. The existence of such an element can be verified by a finite computation, we refer the reader to the remark at the bottom of page 9 of [DR04] for further details.
@@
Owing in part to the fact $(\mathcal{O},\mu)$ has been fixed, there are only two possible values [^3] This follows from [Rot04, Lemmas 3.5 and 3.7]. Indeed, in the notation of the paper, we have $F=\mathbb{Q}$ and $D\neq 3$, as $D$ is divisible by an even number of primes. This forces $\omega_{odd}=1$. The cited lemmas in turn show $C_{2}\cong U_{0}\leq V_{0}\cong C_{2}\times C_{2}$. The non-trivial element of $U_{0}$ can be represented by $\mu=\sqrt{-D}$. This allows us to write $V_{0}=\langle[\mu],[\chi]\rangle$ where the representative $\chi$ may be taken to have reduced norm $m|D$. Moreover, any representative of $[\chi]$ has reduced norm $m$ up to a rational square. Likewise, representatives of the class $[\mu\chi]$ have reduced norm $D/m$ up to rational squares. $\mathrm{n}(\chi)$ can take and the product of these values is equal to $D$. We call $(\mathcal{O},\mu,\chi)$ a twisted principally polarised order, and say $(\mathcal{O},\mu,\chi)$ is of discriminant $D$ and norm $-\mathrm{n}(\chi)$.
@@
As before, we let $L$ denote the endomorphism field of $A$. Dieulefait and Rotger [DR04, Theorem 3.4] showed $\operatorname{{\rm Gal}}(L/K)$ is isomorphic to one of the trivial group, $C_{2}$, or $C_{2}\times C_{2}$. Moreover, they proved in each case $\operatorname{{\rm End}}_{K}^{0}(A)$ is respectively isomorphic to $B$; one of $\mathbb{Q}(\mu),\mathbb{Q}(\chi),\mathbb{Q}(\mu\chi)$; or $\mathbb{Q}$.
This determines (and heavily restricts) the possible images of
$\operatorname{{\rm Gal}}(L/K)\rightarrow\operatorname{{\rm Aut}}(\operatorname{{\rm End}}(A))\cong N_{B^{*}}(\mathcal{O})/\mathbb{Q}^{*}$
@@
is an order of discriminant $D$ in $B$. Moreover, any order $\mathcal{O}^{\prime}$ of discriminant $D$ which contains $\mathbb{Z}[1,mi,j,k]$, satisfies $\mathcal{O}^{\prime}\otimes\mathbb{Z}_{2}=\mathcal{O}\otimes\mathbb{Z}_{2}$.

### v4

@@ 2. Proof of the main theorems
Let $F/K$ be an odd degree abelian extension unramified outside of 2. Then by class field theory, any prime dividing $F\colon K]$ divides either the class number of $K$, or the multiplicative order of the residue field of a prime above 2 in $K$.
@@
As $E/\mathbb{Q}$ is Galois, the reflex field $E^{*}$ is a subfield of $E$ [Shi98, Prop. 28, p. 62]. Moreover, since $A$ is absolutely simple its CM type is primitive. Applying [Shi98, Prop. 30, p. 65] we find the endomorphism field $L$ equals $E^{*}K$. Finally, if $E/\mathbb{Q}$ is an abelian extension, then $E=E^{*}$ by [Shi98, Example (1), p. 63]. ∎
@@
  • $p\in\{7,11,19,43,67,163\}$ and $A$ is isogenous over $\overline{\mathbb{Q}}$ to the power of an elliptic curve with complex multiplication by $\mathbb{Q}(\sqrt{-p})$.
@@
Clearly, we may apply Theorem 2.4. Suppose first $p$ divides $L\colon\mathbb{Q}$]. Then $g\geq 3$ and $A$ is isogenous over $\overline{\mathbb{Q}}$ to the power of an absolutely simple abelian variety with complex multiplication by a proper (non-trivial) subfield $F$ of $\mathbb{Q}(\zeta_{p})$. Thus there is some minimal finite extension $K/\mathbb{Q}$ such that $A_{K}$ contains a simple abelian subvariety $B$ with CM by $F$. The above lemma applies to $B$ and shows $L$, the endomorphism field of $A$, contains $KF$.
By [Goo21, Prop. 3.9] the group $\operatorname{{\rm Gal}}(L/\mathbb{Q})$ is an extension of $C_{p}$ by $C_{2}^{m}$ for some $m$. In particular, any non-trivial element of even order in $\operatorname{{\rm Gal}}(L/\mathbb{Q})$ has either order 2 or $2p$.
Since $p$ does not divide $F\colon\mathbb{Q}$], the cyclic quotient $\operatorname{{\rm Gal}}(F/\mathbb{Q})$ of $\operatorname{{\rm Gal}}(L/\mathbb{Q})$ has order two. Using $F$ is a CM field only ramified at $p$, we find $F=\mathbb{Q}(\sqrt{-p})$ where $p\equiv 3\mod{4}$. In particular $g$ is odd.
We shall now show $F$ has class number one, which by the Baker–Heegner–Stark Theorem will conclude the proof in this case. As $F/\mathbb{Q}$ is a quadratic extension, $B$ is an elliptic curve. Thus by CM theory [Sil94, Thm. II.4.1, p. 121] $KF$ contains $H$ the Hilbert class field of $F$. In particular, $L$ contains $H$, so $H\colon F]$ divides $2^{m}p$ for some $m$. On the other hand, [FG18, Cor. 2.16] tells us every element of $\operatorname{{\rm Gal}}(H/F)$ has order dividing $g$. As $g$ is odd and less than $p$, we find $\operatorname{{\rm Gal}}(H/F)=1$ as claimed.
We now suppose $p$ does not divide $L\colon\mathbb{Q}$] and show $E\neq\mathbb{Q}(\zeta_{p})$. Suppose we had equality, then applying Lemma 2.5 we find $L=\mathbb{Q}(\zeta_{p})$. First let us note that as they are extensions of $\mathbb{Q}$ of coprime degree, we have $L\cap\mathbb{Q}(A[2])=\mathbb{Q}$. If $p=3$, then [Goo21, Thm. 2.9] implies $\operatorname{{\rm End}}(A)$ is a 2-maximal order in $\operatorname{{\rm End}}^{0}(A)\cong\mathbb{Q}(\zeta_{3})$, so Theorem 2.1 gives us that $L\subseteq\mathbb{Q}(A[2])$. But $L\cap\mathbb{Q}(A[2])=\mathbb{Q}$ so $L=\mathbb{Q}$, contradicting the above. Thus we may take $p\geq 5$. Since $L\cap\mathbb{Q}(A[2])=\mathbb{Q}$, [Goo21, Prop. 3.9] implies $\operatorname{{\rm Gal}}(L/\mathbb{Q})$ is an elementary abelian 2-group. However, $\operatorname{{\rm Gal}}(\mathbb{Q}(\zeta_{p})/\mathbb{Q})$ is not an elementary abelian $2$-group, thus we once again have a contradiction. ∎
Let us record some easy corollaries of [Goo21, Thm. 2.9], Theorems 2.1, 2.6 and Lemma 2.5.
@@
Let $K$ be a number field with a real embedding. Let $f\in K[x]$ be an irreducible cubic polynomial, whose discriminant is a square in $K$. Then the elliptic curve defined by $y^{2}=f(x)$ does not have complex multiplication.
Proof.
Let $A$ denote the elliptic curve defined by $y^{2}=f(x)$. The assumptions on $f$ imply $\operatorname{{\rm Gal}}(f)\cong C_{3}$. Let us suppose for a contradiction that $A$ has complex multiplication. As $2$ is a primitive root modulo 3, we may apply [Goo21, Thm. 2.9]. Since $\mathbb{Q}(\zeta_{3})$ does not have any proper CM subfields, the second case of [Goo21, Thm. 2.9] cannot hold. It follows that $(2)$ is inert in $E\coloneqq\operatorname{{\rm End}}^{0}(A)$ and $\operatorname{{\rm End}}(A)$ is 2-maximal in $E$.
By Lemma 2.5, the endomorphism field of $A$ is $KE$. As $E$ does not have any real embeddings, it is not a subfield of $K$. Thus $KE\colon K]=2$. Since $K(A[2])=K(f)$ is an odd degree extension of $K$, it does not contain $KE$. Thus by Theorem 2.1 and the above, we find that $(2)$ is ramified in $E/\mathbb{Q}$. But this contradicts the above, so we find $A$ cannot have complex multiplication as claimed. ∎
@@
Let $A/\mathbb{Q}$ be an abelian surface. Suppose $\operatorname{{\rm Gal}}(\mathbb{Q}(A[2])/\mathbb{Q})\cong C_{5}$. Then either $\operatorname{{\rm End}}(A)\cong\mathbb{Z}$ or $\operatorname{{\rm End}}_{\mathbb{Q}}^{0}(A)=\operatorname{{\rm End}}^{0}(A)\cong\mathbb{Q}(\sqrt{5})$.
Proof.
By Corollary 2.6, we have either $\operatorname{{\rm End}}^{0}(A)\cong\mathbb{Q}$ or $\mathbb{Q}(\sqrt{5})$. In the first case $\operatorname{{\rm End}}(A)\cong\mathbb{Z}$. Thus we may suppose $\operatorname{{\rm End}}^{0}(A)\cong\mathbb{Q}(\sqrt{5})$.
We are left to show that $L$, the endomorphism field of $A$, equals $\mathbb{Q}$. The injection $\operatorname{{\rm Gal}}(L/\mathbb{Q})\hookrightarrow\operatorname{{\rm Aut}}_{\mathbb{Q}}(\operatorname{{\rm End}}^{0}(A))\cong\operatorname{{\rm Gal}}(\mathbb{Q}(\sqrt{5})/\mathbb{Q})$ implies the degree $L\colon\mathbb{Q}$] divides 2. Since $\mathbb{Q}(A[2])=\mathbb{Q}(f)$ is an odd degree extension of $\mathbb{Q}$, we find $L\cap\mathbb{Q}(A[2])=\mathbb{Q}$. Thus $L$ is contained in $\mathbb{Q}(A[2])$ if and only if $L=\mathbb{Q}$.
As 2 is a primitive root modulo 5, we may apply [Goo21, Thm. 2.9] which implies that $\operatorname{{\rm End}}(A)$ is 2-maximal in $\operatorname{{\rm End}}^{0}(A)\cong\mathbb{Q}(\sqrt{5})$. Since 2 is inert in $\mathbb{Q}(\sqrt{5})$, Theorem 2.1 implies that $L\subseteq\mathbb{Q}(A[2])$. Thus $L=\mathbb{Q}$ as claimed. ∎
@@
Both cases are possible. Indeed, examples of the first case are easily found. For the second case, note the jacobian $J$ of the hyperelliptic curve $y^{2}=x(x^{5}-4x^{4}+2x^{3}+5x^{2}-2x-1)$ has $\operatorname{{\rm End}}_{\mathbb{Q}}(J)=\operatorname{{\rm End}}(J)\cong\mathbb{Z}\left[\frac{1+\sqrt{5}}{2}\right]$ by [Wil00, Prop. 1] and the Galois group of $x^{5}-4x^{4}+2x^{3}+5x^{2}-2x-1$ has order 5.
@@
  • $\operatorname{{\rm End}}^{0}(A)$ is isomorphic to a proper subfield of $\mathbb{Q}(\zeta_{p})$; or
  • $p\in\{7,11,19,43,67,163\}$ and $A$ is isogenous over $\overline{K}$ to the power of an elliptic curve with complex multiplication by $\mathbb{Q}(\sqrt{-p})$.
@@
Clearly, we may apply Theorem 2.4. Suppose first $p$ divides $L\colon K]$. Then $A$ is isogenous over $\overline{K}$ to the power of an absolutely simple abelian variety $B$ with complex multiplication by a proper subfield $F$ of $\mathbb{Q}(\zeta_{p})$. Lemma 2.5 applies to $B$ and shows $L$, the endomorphism field of $A$, contains $F$.
By [Goo21, Prop. 3.9] the group $\operatorname{{\rm Gal}}(L/K)$ is an extension of $C_{p}$ by $C_{2}^{m}$ for some $m$. In particular, any non-trivial element of even order in $\operatorname{{\rm Gal}}(L/K)$ has either order 2 or $2p$. Since $p$ does not divide $F\colon\mathbb{Q}$], the cyclic quotient $\operatorname{{\rm Gal}}(FK/K)$ of $\operatorname{{\rm Gal}}(L/K)$ has order dividing two.
In turn we deduce $F\colon\mathbb{Q}$] divides $4$ and is equal to 4 only if $F\supsetneq K$. Being a CM field, $F\colon\mathbb{Q}]=4$ would imply $F$ contains a real quadratic field in addition to the imaginary quadratic field $K$. Thus $\operatorname{{\rm Gal}}(F/\mathbb{Q})\cong C_{2}\times C_{2}$ contradicting the fact $F\subseteq\mathbb{Q}(\zeta_{p})$. Thus $F\colon\mathbb{Q}]=2$. As $F$ is a CM field only ramified at $p$, we deduce $F=\mathbb{Q}(\sqrt{-p})$ where $p\equiv 3\mod{4}$.
Let $H$ be the Hilbert class field of $F=\mathbb{Q}(\sqrt{-p})$. By [FG18, Thm. 2.14] the group $\operatorname{{\rm Gal}}(HK/FK)$ has order dividing $g$, thus arguing as in the proof of Theorem 2.6, we deduce $HK\colon FK]=1$. In other words, $H\subseteq FK=K(\sqrt{-p})$.
@@
We now suppose $p$ does not divide $L\colon K]$ and show $E\neq\mathbb{Q}(\zeta_{p})$. Suppose we had equality, then by Lemma 2.5, $L=K(\zeta_{p})$. Applying [Goo21, Prop. 3.9] shows $\operatorname{{\rm Gal}}(L/K)=\operatorname{{\rm Gal}}(K(\zeta_{p})/K)\cong\operatorname{{\rm Gal}}(\mathbb{Q}(\zeta_{p})/\mathbb{Q}(\zeta_{p})\cap K)$ has order at most $2$. Hence $\mathbb{Q}(\zeta_{p})\colon\mathbb{Q}$] divides $4$, being equal to 4 only if $K$ is contained in $\mathbb{Q}(\zeta_{p})$. But $\mathbb{Q}(\zeta_{p})\colon\mathbb{Q}]=4$ implies $p=5$ and $\mathbb{Q}(\zeta_{5})$ does not have an imaginary subfield. Whence $\mathbb{Q}(\zeta_{p})\colon\mathbb{Q}]=2$ and $p=3$. This in turn implies $g=1$ which we have ruled out by assumption. ∎
@@
Let $K$ be an imaginary quadratic field with class number divisible by 5. Let $H$ be a degree 5 extension of $K$ contained in the Hilbert class field of $K$. What are the possible endomorphism algebras for abelian surfaces $A/\mathbb{Q}$ with $\mathbb{Q}(A[2])=H$?
@@
With reference to the above question, there are already several restrictions on the possible endomorphism algebras. Indeed, by combining Theorem 2.1, [Goo21, Thm. 2.10] and [Shi98, Prop. 30, p. 65; Example (2), p. 64], we see that either $\operatorname{{\rm End}}^{0}(A)\cong\mathbb{Q}$ or $\mathbb{Q}(\sqrt{d})$ with $d\equiv 5\mod{8}$. The above question then becomes: given $H$, which $d$ occur?
We finish this section with the following extension of [Goo21, Thm. 3.5].
@@
Let $f\in K[x]$ be a polynomial of odd degree $n$ with Galois group isomorphic to a Frobenius group $G$ of order n(n1)n(n-1). Let $J_{f}$ be the jacobian associated to the hyperelliptic curve defined by $y^{2}=f(x)$.
Suppose ${\rm End}^{0}(J_{f})$ is isomorphic to a number field $E$ of dimension $s$ over $\mathbb{Q}$. Then $E/\mathbb{Q}$ is Galois with ${\rm Gal}(E/\mathbb{Q})$ isomorphic to a quotient of $H$ the Frobenius complement of $G$.
Furthermore, $L/K$ is an extension of degree $s$ contained in $K(f)$, and as abstract groups, ${\rm Gal}(L/K)\cong{\rm Gal}(E/\mathbb{Q})$. Moreover, if s=n1s=n-1, then $L=EK$.
@@
This result is [Goo21, Thm. 3.5] with the extra assertion that if s=n1s=n-1, then $L=EK$. If s=n1s=n-1, then $J_{f}$ is an absolutely simple abelian variety with CM by $E$. We may therefore apply Lemma 2.5 to find $L=E^{*}K$ and $E^{*}\subseteq E$. Hence $E^{*}\colon\mathbb{Q}]\geq[L\colon K]=s=[E\colon\mathbb{Q}$]. It follows we have equality $E^{*}=E$. ∎
@@
This field is totally inert at 2 (as predicted by [Goo21, Thm. 2.9]) and unramified outside 13. We note its class number is one, but in line with the theorems presented above, $\mathbb{Q}(f)/K$ is a degree 5 extension ramified only at 2.
@@
Let $B=\mathbb{Q}+i\mathbb{Q}+j\mathbb{Q}+k\mathbb{Q}$ be an indefinite quaternion algebra generated by $i,j$ satisfying $i^{2}=D/m$, $j^{2}=m$, ij=jiij=-ji and $k=ij$, where $D$ is a positive squarefree integer and $m|D$. We note that as $B$ is indefinite, $0>-m=\mathrm{n}(j)$, the reduced norm of $j$. As we shall only use the reduced norm, reduced trace and reduced discriminant, we shall refer to these simply as the norm, trace and discriminant respectively.
@@
An order of $B$ is said to be hereditary if it has squarefree discriminant (for alternate, equivalent, definitions see [DR04, Voi21]). We shall take $\mathcal{O}$ to be a hereditary order in $B$ of discriminant $D$.
@@
An abelian surface $A/K$ is said to have quaternion multiplication by $\mathcal{O}$ if there is an isomorphism $\iota\colon\mathcal{O}\xrightarrow{\sim}\operatorname{{\rm End}}(A)$. We shall assume there is an element $\mu\in\mathcal{O}$ satisfying $\mu^{2}+D=0$. This is always the case when $\mathcal{O}$ is a maximal order [Voi21, 43.6.6, p. 818] and furthermore such an element $\mu$ induces a principal polarisation on $A$ [Voi21, p. 821]. Fix such an $A$, $\mu$ and $\iota$.
@@
Following [DR04, Definition 3.3], we say $\chi\in B$ is a twist of $(\mathcal{O},\mu)$ if it lies in both $\mathcal{O}$ and the normaliser $N_{B^{*}}(\mathcal{O})$, has trace zero, $\mathrm{n}(\chi)$ divides $D$, and $\mu\chi=-\chi\mu$. The existence of such an element can be verified by a finite computation, we refer the reader to [DR04, Rem., p. 9] for further details.
@@
Owing in part to the fact $(\mathcal{O},\mu)$ has been fixed, there are only two possible values [^2] This follows from [Rot04, Lemmas 3.5 and 3.7]. Indeed, in the notation of the paper, we have $F=\mathbb{Q}$ and $D\neq 3$, as $D$ is divisible by an even number of primes. This forces $\omega_{odd}=1$. The cited lemmas in turn show $C_{2}\cong U_{0}\leq V_{0}\cong C_{2}\times C_{2}$. The non-trivial element of $U_{0}$ can be represented by $\mu=\sqrt{-D}$. This allows us to write $V_{0}=\langle[\mu],[\chi]\rangle$ where the representative $\chi$ may be taken to have reduced norm $m|D$. Moreover, any representative of $\chi]$ has reduced norm $m$ up to a rational square. Likewise, representatives of the class $\mu\chi]$ have reduced norm $D/m$ up to rational squares. $\mathrm{n}(\chi)$ can take and the product of these values is equal to $D$. We call $(\mathcal{O},\mu,\chi)$ a twisted principally polarised order, and say $(\mathcal{O},\mu,\chi)$ is of discriminant $D$ and norm $\mathrm{n}(\chi)$.
@@
As before, we let $L$ denote the endomorphism field of $A$. Dieulefait and Rotger [DR04, Thm. 3.4] showed $\operatorname{{\rm Gal}}(L/K)$ is isomorphic to one of the trivial group, $C_{2}$, or $C_{2}\times C_{2}$. Moreover, they proved in each case $\operatorname{{\rm End}}_{K}^{0}(A)$ is respectively isomorphic to $B$; one of $\mathbb{Q}(\mu),\mathbb{Q}(\chi),\mathbb{Q}(\mu\chi)$; or $\mathbb{Q}$. This determines (and heavily restricts) the possible images of
$\operatorname{{\rm Gal}}(L/K)\hookrightarrow\operatorname{{\rm Aut}}_{\mathbb{Z}}(\operatorname{{\rm End}}(A))\cong N_{B^{*}}(\mathcal{O})/\mathbb{Q}^{*}$
@@
is an order of discriminant $D$ in $B$. Moreover, any order $\mathcal{O}^{\prime}$ of discriminant $D$ which contains $\mathbb{Z}[1,mi,j,k]$, satisfies $\mathcal{O}^{\prime}\otimes\mathbb{Z}_{(2)}=\mathcal{O}\otimes\mathbb{Z}_{(2)}$.